OpenStudy (anonymous):

If cos = -2/3, what do csc,tan, and sin equal?

OpenStudy (anonymous):

use \[\cos ^{2}x + \sin ^{2}x = 1\] this will give u sinx. then u can find tan and cosec using the trigo ratios.

OpenStudy (anonymous):

do i put -2/3 in place of x? and what are the trigo ratios?

OpenStudy (anonymous):

-2/3 is the value given for cosx, not just x. so substitute cosx = -2/3 in the above equation to get (-2/3)^2 + sin^2 x = 1. this will give u the value of sinx. then tanx = sinx / cosx and cosecx = 1 / sinx.